### v1

# DLinear-based Prediction of Remaining Useful Life of Lithium-Ion Batteries: Feature Engineering through Explainable Artificial Intelligence

Abstract
Accurate prediction of the Remaining Useful Life (RUL) of lithium-ion batteries is essential for ensuring safety, reducing maintenance costs, and optimizing usage. However, predicting RUL is challenging due to the nonlinear characteristics of the degradation caused by complex chemical reactions. Machine learning allows precise predictions by learning the latent functions of degradation relationships based on cycling behavior. This study introduces an accurate RUL prediction approach based on feature engineering and DLinear, applied to the dataset from NASA’s Prognostics Center of Excellence. Among the 20 features generated from current, voltage, temperature, and time provided in this dataset, key features contributing to degradation are selected using Pearson correlation coefficient and Shapley values. Shapley value-based feature selection effectively reflects cell-to-cell variability, showing similar importance rankings across all cells. The DLinear-based RUL prediction using key features efficiently captures the time-series trend, demonstrating significantly better performance compared to Long Short-Term Memory and Transformer models.
11footnotetext: These authors contributed equally to this work.
Keywords— Lithium-Ion Batteries, Remaining Useful Life, Feature Engineering, DLinear, Cell-to-Cell Variability
1 Introduction
The Electric Vehicles (EVs) industry has grown rapidly due to environmental issues related to internal combustion engine vehicles. Lithium-Ion Batteries (LIBs) are the most widely used battery materials based on their advantages such as high energy density, low self-discharge characteristics, and long lifespan [1]. However, it is inevitable that battery performance will degrade due to capacity fade and impedance increase that occurs during cycling. Monitoring capacity reduction and accurately predicting Remaining Useful Life (RUL) through battery management systems have become crucial, as the performance degradation of lithium-ion batteries can end up with critical risks such as thermal runaway reactions [2, 3]. Predicting the necessary replacement time in advance is challenging, due to intra-cell factors including morphology, architecture, and composition, as well as inter-cell factors such as winding, connection, and assembly processes that contribute to Cell-to-Cell Variability (CtCV) [4, 5], along with differences in operating conditions for each device. The Prognostics and Health Management (PHM) system implements condition-based management by predicting the necessary replacement timing before failures occur, allowing for repairment only when it’s necessary [6, 7]. It considers battery capacity degradation using various sensor data, such as voltage, current, and temperature. RUL prediction is one of the critical functions of the PHM system, which forecasts the remaining cycles where the battery can be used safely. Accurate RUL prediction contributes to optimizing battery replacement timing and prevents accidents caused by excessive charging or discharging.
RUL prediction methods using early cycling behavior can generally be classified into two approaches: model-based approach and data-driven approach [8, 9]. However, model-based approaches often encounter challenges. In practice, the irreversible capacity fade is caused by an interaction of mechanical (e.g., particle cracking) and chemical (e,g., electrolyte decomposition) factors, making it quite difficult to mathematically identify these degradation mechanisms. Additionally, there can be a knee point where performance degradation, which progresses slowly in the early stages, accelerates after a certain cycle. This nonlinearity of performance degradation also contributes to the difficulty of model-based approaches. Machine Learning (ML)–based approach is a promising alternative to overcome this challenge [10, 11, 12].
ML–based RUL prediction research typically employs time series algorithms, with changes observed for each charging/discharging cycle. Traditionally, Recurrent Neural Network (RNN) [13, 14, 15] and Long Short-Term Memory (LSTM) [16] have been used for such tasks: Liu et al. [17] propose an adaptive RNN that continuously updates the parameters of the network to improve the description performance of dynamic cycling behavior of the LIBs. Park et al. [18] conducted RUL forecasting using LSTM, but they struggle with learning long-term dependency [19]. To overcome this limitation, the attention algorithm was introduced, enabling predictions that consider a broader context rather than just neighboring time steps [20, 21]. Designed for natural language processing task and centered around the attention algorithm, Transformer architectures outperforms RNN and LSTM by taking a bird’s-eye view of time-series data to learn features of data, then trains in order of importance rather than sequential dependency [22]. Chen et al. [23] conducted a prediction of RUL by Transformer, successfully overcoming the long-term dependency hurdle of RNN and LSTM. Despite being a game changer in time series forecasting, the Transformer has encountered difficulties when applied to time series prediction. Specifically, its high complexity in space (i.e., memory) time and its performance are not significantly superior to conventional LSTM models.
An alternative, DLinear, was proposed to address these challenges [24]. DLinear uses two linear layers considering trend and seasonality, in contrast to the complex structure and extensive computation of Transformer. Unlike the Transformer, DLinear takes both trend and seasonality into account in time series data predictions. Furthermore, it provides explainablity with its weights of linear layers, revealing parameter importance. For battery life, RUL typically follows a decreasing trend with occasional increases. While the Transformer struggles to learn this trend, DLinear captures both the decreasing and periodic increasing trends, outperforming LSTM and Transformer-based predictions in aspects of space complexity and time complexity.
The performance of ML–based RUL prediction such as DLinear is determined by the quantity and quality of data. Since battery cycling experiments require a significant budget in time and cost, there are practical difficulties in generating large-scale datasets. Therefore, it is important to identify features that significantly contribute to battery lifespan and achieve high prediction performance even with a small number of training cycles. However, because high-dimensional features can overfit the cycling behavior used for training, selecting key features is essential, and various feature engineering studies have been conducted to predict battery lifespan: Fei et al. [25] extracted 42 features from a large-scale lithium iron phosphate (LFP) battery dataset [11] and identified a high-performance prediction model structure by applying various feature selection approaches such as wrapper and filter approaches and various prediction models. Hu et al. [26] proposed a fusion approach that combines the advantages of wrapper and filter approaches to select features that contribute to state-of-health prediction. Greenbank et al. [27] propose an automatic feature selection method using the Pearson Correlation Coefficient (PCC), a filter approach, to predict the knee point where battery life declines rapidly and input the End-of-Life (EoL) prediction. However, the most widely used filter approach has concerns that correlation ranking can only consider linear dependencies between features and output. Shapley Additive exPlanation (SHAP) [28], which has recently been used in sensitivity analysis, provides insight at the feature selection stage because it presents qualitative results as well as quantitative sensitivity [29, 30]. Additionally, in Ref. [31], SHAP value-assisted feature selection was found to have superior performance compared to other algorithms for time series applications such as battery lifespan prediction. Moreover, the SHAP-based approach, which is a model-agnostic approach, is useful when applied to various machine learning models. In this study, we identify features that contribute to RUL prediction through the filter approach and SHAP-based approach and apply them to a ML model.
Here, we identify the features that contribute to the battery lifespan and predict the RUL of the battery using the DLinear model. Features contributing to battery capacity degradation were extracted based on battery chemistry. For the entire feature set, each key feature subset identified through the filter approach and SHAP-based approach is used as input for LSTM, Transformer, and DLinear, and the accuracy and EoL cycle are compared. The rest of this paper is structured as follows. In Section 2, the three machine learning algorithms used for the prediction of RUL are described. In Section 3, the dataset used in the cycling experiment and the scheme for extracting features contributing to the battery lifespan based on battery chemistry are explained. Additionally, two algorithms used for feature selection are introduced. Section 4 discusses feature engineering and RUL prediction results, and Section 5 presents a summary of the results and future work.
2 Time-series prediction model
Figure 1: Schematic diagram of RUL prediction through feature engineering
2.1 Long short-term memory (LSTM)
Sequential Neural Network (SNN) process time-series data by assuming that the state at any given time step depends on prior time steps. RNN is the pioneering SNN structure designed for processing time-series data [13, 14, 15]. The RNN model takes input to produce output, and also sends its output back to itself, to recurrent into the next time step. Due to this structure, each neuron in RNN has two sets of weights: one for input and another for multiplying with the output from the previous time step. In this way, RNN takes previous information into account and determines the extent of this reference with weights. However, it has pointed that as time progresses, the continuous multiplication of weights in RNN results in the dilution of information from distant past time steps at the current time step, a problem known as gradient vanishing [32].
To address this issue, the SNN known as LSTM was suggested, which divides the output from the previous time step into long-term state and short-term state. LSTM forwards them to the subsequent time steps. By introducing long-term state with minimal loss according to weights, LSTM improves the gradient vanishing problem observed in RNN.
2.2 Transformer
For a long time, LSTM has been firmly established as the state-of-the-art approach in time series data. However, long-term state of LSTM also did not completely solve the gradient vanishing problem [19]. The problem was alleviated but similar, as the consideration of information from distant past time steps being diluted as weights were multiplied with the long-term state of LSTM with the progression of time steps. Over time, an Attention algorithm was devised to focus more on contextually significant information after receiving information, not only from the immediate previous time step but from all time steps [20, 21]. This allowed for learning based on meaningful information from the entire dataset rather than focusing solely on nearby information, thereby addressing the issues of RNN, LSTM, and similar SNN algorithms. Furthermore, implemented with the Attention algorithm, the Transformer was devised. Recently, Transformer-based approaches and variant models of Transformer were utilized in predicting the RUL of LIBs [23].
2.3 DLinear
The Transformer has successfully overcome the hurdle of gradient vanishing and long-range dependencies posed by LSTM. However, as the Transformer was designed for natural language processing tasks (e.g., translation, summarization), not for time-series data, it had the weakness of poor performance when applied to time-series data. In particular, early Transformer models had difficulty in identifying trends in data that increased or decreased continuously throughout the observation period. Furthermore, due to its complex structure for analyzing sentences and texts written in languages, the Transformer required excessive computation compared to the conventional SNNs, thereby demanding a significant amount of time and memory. To address these two issues, DLinear was proposed [24].
Algorithm 1 Decomposition-Linear (DLinear) Model for RUL Prediction
1:Input: Feature matrix $X:[F_{1},F_{2},...,F_{20}]$, Target $y=\text{RUL}$
2:Step 1: Compute the moving average $X_{t}$ using a sliding window
3:Step 2: Decompose into trend ($X_{t}$) and remainder ($X_{s}$) with $X_{s}=X-X_{t}$
4:Step 3: Pass $X_{t}$ and $X_{s}$ through linear layers
5:if $individual$ is True then
6:    for each feature $i$ in $X$ do
7:         Apply linear layers to predict future trend and seasonal components
8:    end for
9:else
10:    Use shared linear layers for all features
11:end if
12:Step 4: Combine predicted $X_{t}$ and $X_{s}$ to get final prediction
13:Output: Predicted time series data
The detailed procedure for RUL prediction using DLinear is introduced in Algorithm 1. DLinear separates time-series data into Trend ($X_{t}$) and Remainder ($X_{s}$) using a moving average, and learns each through a separate linear layer. Equation of $X_{t}$ and $X_{s}$ can be defined as follows:
$X_{t}=\frac{1}{n}\sum_{k=0}^{n-1}X_{i-k}$ (1)
$X_{s}=X-X_{t}$ (2)
where $i$ is the time step of data, and $n$ is the window size of the moving average. Based on the learning of these two linear layers, the future Trend and Remainder are predicted, and these two predictions are combined to derive the final prediction.
3 Feature engineering
In this section, feature engineering to extract parameters contributing to RUL from cycling data of LIBs and identify key parameters is explained. Our RUL prediction framework is validated using the NASA dataset [33]. In Section 3.1, the process of extracting features expected to contribute to lifespan through information on voltage, temperature, and time for four LIBs cells (i.e., B0005, B0006, B0007, and B0018) is described. In Section 3.2, the process by which a significantly contributing subset of the entire feature set is selected through the filter approach and SHAP-based approach is explained.
3.1 Feature extraction
In feature extraction, factors that contribute to RUL prediction are extracted based on battery cycling data collected from the Battery Management System (BMS). According to battery chemistry, the main factors that accelerate degradation are voltage [34] and temperature [35] during cycling, and degradation causes differences in charge and discharge profiles. Table 1 describes 20 features categorized by current and voltage, temperature, and time. Fig. 2(a-c) shows that the current and voltage profile at the charging stage, temperature distribution, and discharge voltage curve are dependent on the RUL. These dependent characteristics are processed through statistical information such as skewness and included in the feature set. In particular, the characteristic of the discharge curve, where differences occur after significant cycles, is explained by the slope at the early stage of the discharge (Fig. 2(d)).
Table 1:
Description of features extracted for each parameter
Related parameter Feature designation Description
Current and Voltage F1 Variance of measured voltage in discharging
F2 Variance of measured current in discharging
F3 Median of loaded voltage in discharging
F4 Skewness of measured voltage in discharging
F5 Skewness of loaded voltage in discharging
F6 Slope of discharge voltage curve (50s to 500s)
F7 Slope of discharge voltage curve (50s to 1000s)
F8 Slope of discharge voltage curve (50s to 1500s)
Temperature F9 Maximum temperature in discharging
F10 Average temperature in discharging
F11 Variance of temperature in discharging
F12 Skewness of temperature in discharging
F13 Minimum temperature in discharging
F14 Maximum temperature in charging
F15 Minimum temperature in charging
F16 Average temperature in charging
F17 Skewness of temperature in charging
Time F18 CC charging time
F19 CV charging time
F20 Total discharging time
Figure 2: Feature description of charging and discharging for B0005 battery cell: (a) CC charging time (F18), (b) Temperature in discharging, (c) Voltage in discharging, (d): Slope of discharge voltage curve (F6, F7, and F8)
3.2 Feature selection
3.2.1 Filter approach
The filter approach uses feature ranking to select key features. In general, the correlation between the extracted features and the objective function (i.e., RUL) scores the ranking of each feature, selects features that are higher than a certain threshold, and screens other features [36]. Because the filter approach is applied without being integrated into the prediction model, it can be easily applied to high-dimensional feature sets with relatively low computational budgets. This quantifies the linear correlation between the feature denoted by $x$ and the objective function denoted by $y$.
$r_{xy}=\frac{\sum_{i=1}^{n}(x_{i}-\bar{x})(y_{i}-\bar{y})}{\sqrt{\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}}.$ (3)
Where $\bar{x}$ and $\bar{y}$ are sample mean of $x$ and $y$, respectively. As for the key feature selection, all correlation coefficients between key features and cell capacity are evaluated based on Eq. 3. The three parameter subsets with the highest absolute correlation coefficients are regarded as the most sensitive features, while those with the lowest absolute values are considered the least sensitive features. This concept also applies to the SHAP-based approach described in Section 3.2.2
3.2.2 SHAP-based approach
SHAP is an approach derived from Shapley values to explain the contribution of each feature to the predictions of a machine learning model. Originating from game theory, Shapley values aim to fairly distribute the contribution when players collectively achieve a specific outcome. Owen et al. [29] identified the relationship between Shapley values and variance-based global sensitivity analysis. The sum of the Shapley values for all features used in training a machine learning model is exactly equal to the total variance. This normalization property is quite crucial when determining the importance ranking of features. Sobol’ indices, widely used in global sensitivity analysis, provide contributions for individual features similarly to SHAP, but when dependencies exist between features, the sum of sensitivities may not match the total variance [37]. In contrast, Shapley values work well even in the presence of dependencies between features. The Shapley value for feature Fi of a machine learning model is as follows:
$\phi_{i}=\sum_{S\subseteq N\setminus\left\{i\right\}}\frac{\left|S\right|!\left(\left|N\right|-\left|S\right|-1\right)!}{\left|N\right|!}\left[f\left(S\cup\left\{i\right\}\right)-f\left(S\right)\right].$ (4)
Where $\phi_{i}$ is the Shapley value of the $i$-th feature Fi, $N$ is the set of all features (e.g., in our case, $\left|N\right|$ is 20, see the Table 1), $N\setminus\left\{i\right\}$ is the set of all possible combinations of remaining features excluding feature Fi, $S$ is a subset of $N\setminus\left\{i\right\}$, $f\left(S\cup\left\{i\right\}\right)$ is the model prediction of the union of Fi and feature set $S$, and $f$($S$) is the model prediction for the features in $S$.
4 Results and discussion
Figure 3: PCC of all parameters and RUL for four battery cells: (a) B0005 cell, (b) B0006 cell, (c) B0007 cell, (d) B0018 cell
Figure 4: SHAP score of all parameters and RUL for four battery cells: (a) B0005 cell, (b) B0006 cell, (c) B0007 cell, (d) B0018 cell
The current, voltage, temperature, and time data collected from the cycling experiment were reproduced into 20 features as shown in Table 1. Fig. 3 presents the results of PCC on 20 features for four cells. The three most sensitive features for each cell, as identified by the PCC, are:
  • B0005: F10 (Average temperature in discharging), F12 (Skewness of temperature in discharging), F13 (Minimum temperature in discharging)
  • B0006: F3 (Median of loaded voltage in discharging), F18 (CC charging time), F20 (Total discharging time)
  • B0007: F7 (Slope of discharge voltage curve, 50s to 500s), F8(Slope of discharge voltage curve, 50s to 1000s), F18(CC charging time)
  • B0018: F7 (Slope of discharge voltage curve, 50s to 500s), F8 (Slope of discharge voltage curve, 50s to 1000s), F20 (Total discharging time)
B0006, B0007, and B0018 include voltage characteristics during the discharging stage and time-related parameters during the charging or discharging stages as key parameters. In contrast, all three key parameters for B0005 are related to temperature characteristics, demonstrating significant variability across the four cells.
Fig. 4 shows the results of SHAP analysis, that F2 (Variance of measured voltage in discharging) and F11 (Variance of temperature in discharging) were selected as the highest values across all four cells. In addition to F2 and F11, F10 (Average temperature in discharging) was identified as one of the three most sensitive feature for B0005. Meanwhile, F18 (CC charging time), along with F2 and F11, was identified as one of the three most sensitive feature for B0006, B0007, and B0018.
When training with the existing 70 cycles, the EoL prediction gap was well captured in Fig. 5. EoL was defined as 80% of the nominal capacity. In the case of LSTM, it captured and tracked the decreasing trend well after the training cycles. However, as the prediction progressed toward the later stages, it failed to follow the downward trend, resulting in a flat graph. This is because of the gradient vanishing problem mentioned previously.
The performance of Transformer was found to be inferior than LSTM due to the issue of over-stationarization, which is considered a drawback of the Transformer algorithm [38]. Since the Transformer was originally developed for natural language processing rather than time-series data, it fails to capture critical temporal dependencies in non-stationary time-series data with clear downward trends. Despite the fact that forecasting non-stationary time-series data is crucial for real-world applications including battery RUL predictions, such over-stationarization limits the prediction capability of the Transformer in these non-stationary time-series data. Previous studies on RUL prediction using Transformers [23, 39] have demonstrated that Transformers outperform LSTM in terms of predictive performance with revised model of Transformer. However, this study focuses on a comprehensive analysis of feature engineering considering CtCV and the improvement in RUL prediction performance driven by the characteristics of the DLinear algorithm. In this context, optimization of the model architecture was not considered in this study.
While LSTM and Transformer show a significant decrease in prediction performance as the training period decreases, DLinear maintained relatively accurate EoL prediction performance. Even when the training cycles were reduced to a maximum of 30 cycles, The performance of DLinear was observed to decline. Thus, DLinear demonstrated excellent performance in battery RUL prediction by separating the trend and seasonality of time-series data and learning them linearly, thereby effectively overcoming both the gradient vanishing problem of LSTM and the over-stationarization problem of the Transformer.
Figure 5: RUL prediction of four LIBs : (a) B0005, (b) B0006, (c) B0007, (d): B0018
Figure 6: Comparison of R2 score according to algorithm and training cycle
$R^{2}=1-\frac{SS_{\text{res}}}{SS_{\text{tot}}}$ (5)
Fig. 6 illustrates the R2 scores of the models based on each algorithm, while Table 2 presents the MSE values according to the algorithm and the size of the training cycles, and Table 3 shows the corresponding R2 scores. The R2 score is not affected by the scale of the data, therefore particularly useful for evaluating battery data with varying training cycle lengths and total cycle durations. In contrast, MSE, by squaring the errors, sensitively captures large errors in time-series data and is differentiable, making it a commonly used loss function in machine learning algorithms. When evaluated using MSE and R2 score, DLinear’s performance was outstanding. The bolded values in Table 2 and Table 3 indicate the superior RUL prediction performance of DLinear. Even when the training cycles were reduced to 40, DLinear achieved an R2 score of 0.995. $SS_{\text{res}}$ is the sum of squared differences between the observed values and the predicted values, and $SS_{\text{tot}}$ is the sum of squared differences between the observed values and the mean of the observed data. In cases where the model performs worse than simply predicting the mean of the observed values, the R2 score will be negative. This is because Eq. 5 subtracts the ratio of the residual sum of squares from 1, meaning the resulting negative value is not constrained within the range of -1 to 1.
Figure 7: RUL prediction of B0005 according to training cycle : (a) 70 cycle, (b) 50 cycle, (c) 40 cycle, (d): 30 cycle
Figure 8: RUL prediction of B0006 according to training cycle : (a) 70 cycle, (b) 50 cycle, (c) 40 cycle, (d): 30 cycle
Figure 9: RUL prediction of B0007 according to training cycle : (a) 70 cycle, (b) 50 cycle, (c) 40 cycle, (d): 30 cycle
Figure 10: RUL prediction of B0018 according to training cycle : (a) 70 cycle, (b) 50 cycle, (c) 40 cycle, (d): 30 cycle
Table 2: Comparison of MSE by algorithm according to training cycle
Algorithm Training Cycle
70 50 40 30
B0005 LSTM 0.03407 0.01821 0.01738 0.01999
Transformer 0.15731 0.31750 0.11027 0.37554
DLinear 0.00003 0.00009 0.00019 0.00042
B0006 LSTM 0.02038 0.00924 0.01089 0.01532
Transformer 0.03699 0.24336 0.60980 0.35563
DLinear 0.00003 0.00006 0.00009 0.00019
B0007 LSTM 0.04070 0.01757 0.01530 0.01355
Transformer 0.02553 0.19942 0.10073 0.15515
DLinear 0.00007 0.00031 0.00062 0.00220
B0018 LSTM 0.01101 0.03776 0.02583 0.01577
Transformer 0.02165 0.02092 0.07549 0.20434
DLinear 0.00002 0.00003 0.00003 0.00006
Table 3: Comparison of R2 by algorithm according to training cycle
Algorithm Training Cycle
70 50 40 30
B0005 LSTM -0.17398 0.66193 0.74823 0.75999
Transformer -2.10452 -3.59821 -0.59773 -1.68900
DLinear 0.98387 0.99705 0.99628 0.68536
B0006 LSTM -0.37580 0.63463 0.69113 0.66219
Transformer -1.49730 8.62568 16.29935 6.84430
DLinear 0.99772 0.34021 -0.73222 -0.96263
B0007 LSTM -0.69123 0.61148 0.74657 0.81998
Transformer -0.06079 3.40917 -0.66853 1.06123
DLinear 0.97061 0.99422 0.99511 0.33864
B0018 LSTM -0.57772 -0.40026 0.38087 0.67313
Transformer -2.10273 0.22419 -0.80949 3.21659
DLinear 0.99418 0.98418 0.99464 0.95958
In PHM technologies, the ultimate goal is to achieve accurate predictions with fewer training cycles as possible. To evaluate whether this study can contribute to PHM technology, we progressively reduced the training cycles to 70, 50, 40, and 30 and analyzed the prediction performance. Fig. 7, 8, 9 and, 10 present the prediction results for four different battery cells (B0005, B0006, B0007, and B0018, respectively) while the number of training cycles was reduced. The corresponding MSE and R2 values are shown in Table 2 and Table 3. Reducing training cycles, LSTM exhibited poor performance, with its predictions deviating significantly from the actual values in the later stages, as shown in Fig. 5. Similarly, due to the over-stationarization issue, predictions made using the Transformer also demonstrated severe fluctuations which was to minimize the loss function, consistent with the behavior observed in Fig. 5. On the other hand, except for B0006 cell (Fig. 8), DLinear demonstrates superior prediction performance compared to LSTM and Transformer even at 50 cycles and 40 cycles. However, except for B0018 (Fig. 10), there is a considerable difference between the real cycling behavior and the predicted cycling behavior when the training cycles are set to 30. In fact, recent research using Transformer demonstrates superior predictive performance compared to LSTM [40, 23]. However, this study focuses on feature engineering, and by using a simple network without considering improvements to the machine learning algorithms or hyperparameter tuning, it makes it easier to reproduce the results or apply this framework to other battery systems.
SHAP and PCC were employed for feature selection. Although using the three most sensitive features of performance of PCC resulted in higher prediction performance, there was a lack of consistency across different batteries (B0005, B0006, B0007, B0018, respectively). In contrast, while the three most sensitive features identified using SHAP yielded slightly lower prediction performance than the performance of PCC, they demonstrated consistency in feature selection across different batteries. This result suggests that SHAP-based feature selection has a higher potential for transfer learning between batteries, suggesting a certain level of robustness. In battery capacity research, sampling is often conducted to ensure representativeness that accounts for differences between cells, underscoring the importance of generalization performance. The consistent selection of the same features across multiple batteries indicates that these features are likely useful in a general context rather than being specific to individual cells. This consistency also implies potential reductions in future data collection and analysis costs, making it significant for broader applications.
Figure 11: RUL prediction of B0005 cell according to feature selection: (a) PCC, (b) SHAP
Table 4: The most sensitive and least sensitive features in terms of importance, as determined by SHAP analysis
B0005 B0006 B0007 B0018
Three most sensitive 2a,*, 10, 11a 2a, 11a, 18a 2a, 11a, 18a 2a, 11a, 18a
Three least sensitive 1, 7, 20 1, 5b, 14b 5b, 14b, 16 5b, 14b, 16
* Based on the feature selection results using SHAP, the extracted features, denoted by superscript a, are identified as key features in three or more cells, while those, denoted by superscript b, are identified as not significant in three or more cells.
Table 5: The three most sensitive and three least sensitive features in terms of importance, as determined by PCC
B0005 B0006 B0007 B0018
Three most sensitive 10, 13, 12 18, 20, 3 7, 8, 18 7, 8, 20
Three least sensitive 17, 19, 20 13, 17, 15 13, 14, 15 10, 14, 16
Fig. 11 show the prediction results for B0005 cell using the sets of the three most sensitive features and the three least sensitive features, as analyzed through PCC and SHAP, respectively. The sets of the three most sensitive parameters and the three least sensitive parameters, as determined by SHAP and PCC analyses, are presented in Table 4 and Table 5.
SHAP analysis reveals that more than half of the three most sensitive features are shared among batteries. A noteworthy point is, the key features identified by PCC differ significantly across the four cells, whereas those identified by SHAP analysis are identical for B0006, B0007, and B0018, wich B0005 sharing two features with the other cells. A possible interpretation of these results is that the key parameters identified through SHAP-based sensitivity analysis provide robust results concerning CtCV. In practice, the uncertainties propagated by various causes, such as differences in electrode thickness during manufacturing stages (e.g., Calendaring process), imply that cycling behavior can vary even when cycling the same battery system under identical conditions [5]. The high robustness of SHAP suggests that it is a promising approach for predicting the RUL of large-capacity batteries (e.g., EVs) used in modules or packs.
Meanwhile, as shown in Table 5, PCC revealed cases where a feature ranked low for one battery was ranked high for another. For example, F10, F13, and F20 were selected as the three most sensitive features based on PCC for some batteries, while they were in the three least sensitive features for others. These three features are related to temperature and time, and since batteries are significantly influenced by their surroundings, especially the ambient temperature, measurements such as maximum or minimum values can be affected by external factors or measurement errors. Therefore, these features appear to lack consistency in their significance across different batteries when evaluated based on PCC.
Figure 12: RUL prediction of B0005 cell using two most sensitive features identified by SHAP: F2, F11
In Table 4, F2 and F11 were consistently selected as the three most sensitive features across all four batteries. Building on this observation, we also examined the performance of predictions using only these two features. However, the performance did not match that of predictions made using three features. This suggests that, in addition to F2 and F11, including additional features such as F18, which encapsulate lifespan-related information, is recommended for improved prediction accuracy.
5 Conclusion
PHM technology involves real-time diagnosis and prediction of the health state of a target system. Accurate RUL prediction helps implement condition-based management, where interventions are taken only when necessary by predicting the preemptive system replacement point before failures or defects occur. However, predicting failures through RUL requires measuring charge/discharge data during real usage, which is impractical and time-consuming during experiments. Therefore, it’s crucial to predict with higher accuracy using fewer experiments. In cases of clear trend-based predictions like battery degradation, DLinear outperforms models utilizing other algorithms. Even when trained on the same number of cycles, DLinear can predict further and more accurately than conventional LSTM or Transformer models. This suggests its potential contribution to implementing Condition Based Management for battery PHM technology in real industrial applications. On the other hand, identifying key features for predicting RUL aids in the implementation of PHM technology. PCC yields good performance for individual cells, but the key features vary significantly due to CtCV. Additionally, SHAP does not perform as well as PCC in terms of overall performance. However, SHAP consistently identifies key features, making it advantageous for battery research requiring generalization, despite the CtCV. Particularly, CtCV arises from both intra-cell factors such as morphology, architecture, and composition, and inter-cell factors like winding, connection, and assembly processes. Therefore, creating homogeneous batteries is challenging. In this regard, it is noteworthy that SHAP evaluates the importance of the same features despite CtCV.
Code Availability
The code used in this work will be available after publication.

### v2

@@
-DLinear-based Prediction of Remaining Useful Life of Lithium-Ion Batteries: Feature Engineering through Explainable Artificial Intelligence
+State-of-Health Prediction for EV Lithium-Ion Batteries via DLinear and Robust Explainable Feature Selection
 Abstract
-Accurate prediction of the Remaining Useful Life (RUL) of lithium-ion batteries is essential for ensuring safety, reducing maintenance costs, and optimizing usage.
-However, predicting RUL is challenging due to the nonlinear characteristics of the degradation caused by complex chemical reactions.
-Machine learning allows precise predictions by learning the latent functions of degradation relationships based on cycling behavior.
-This study introduces an accurate RUL prediction approach based on feature engineering and DLinear, applied to the dataset from NASA’s Prognostics Center of Excellence.
-Among the 20 features generated from current, voltage, temperature, and time provided in this dataset, key features contributing to degradation are selected using Pearson correlation coefficient and Shapley values.
-Shapley value-based feature selection effectively reflects cell-to-cell variability, showing similar importance rankings across all cells.
-The DLinear-based RUL prediction using key features efficiently captures the time-series trend, demonstrating significantly better performance compared to Long Short-Term Memory and Transformer models.
+Accurate prediction of the state-of-health (SOH) of lithium-ion batteries is essential for ensuring the safety, reliability, and efficient operation of electric vehicles (EVs).
+Battery packs in EVs experience nonuniform degradation due to cell-to-cell variability (CtCV), posing a major challenge for real-time battery management.
+In this work, we propose an explainable, data-driven SOH prediction framework tailored for EV battery management systems (BMS).
+The approach combines robust feature engineering with a DLinear.
+Using NASA’s battery aging dataset, we extract twenty meaningful features from voltage, current, temperature, and time profiles, and select key features using Pearson correlation and Shapley additive explanations (SHAP).
+The SHAP-based selection yields consistent feature importance across multiple cells, effectively capturing CtCV.
+The DLinear algorithm outperforms long short-term memory (LSTM) and Transformer architectures in prediction accuracy, while requiring fewer training cycles and lower computational cost.
+This work offers a scalable and interpretable framework for SOH forecasting, enabling practical implementation in EV BMS and promoting safer, more efficient electric mobility.
 11footnotetext: These authors contributed equally to this work.
-Keywords— Lithium-Ion Batteries, Remaining Useful Life, Feature Engineering, DLinear, Cell-to-Cell Variability
+Keywords— Lithium-ion batteries, State-of-health, Feature engineering, DLinear, Cell-to-cell variability
 1 Introduction
 The electric vehicles (EVs) industry has grown rapidly due to environmental issues related to internal combustion engine vehicles.
-Lithium-Ion Batteries (LIBs) are the most widely used battery materials based on their advantages such as high energy density, low self-discharge characteristics, and long lifespan [1].
-However, it is inevitable that battery performance will degrade due to capacity fade and impedance increase that occurs during cycling.
-Monitoring capacity reduction and accurately predicting Remaining Useful Life (RUL) through battery management systems have become crucial, as the performance degradation of lithium-ion batteries can end up with critical risks such as thermal runaway reactions [2, 3].
+Lithium-ion batteries (LIBs) are the most widely used battery materials based on their advantages such as high energy density, low self-discharge characteristics, and long lifespan [1, 2, 3].
+However, it is inevitable that battery performance degrades as impedance increases due to chemical decomposition and lithium plating during battery usage.
+Monitoring capacity reduction and accurately predicting state-of-health (SOH) through battery management system (BMS) have become crucial, as the performance degradation of LIBs can pose critical risks, such as thermal runaway reactions [4, 5, 6].
 Predicting the necessary replacement time in advance is challenging, due to intra-cell factors including morphology, architecture, and composition, as well as inter-cell factors such as winding, connection, and assembly processes that contribute to cell-to-cell variability (CtCV) [7, 8], along with differences in operating conditions for each device.
-The Prognostics and Health Management (PHM) system implements condition-based management by predicting the necessary replacement timing before failures occur, allowing for repairment only when it’s necessary [6, 7].
+The prognostics and health management (PHM) system implements condition-based management by predicting the necessary replacement chance before failures occur, allowing for repairment only when it’s necessary [9, 10].
 It considers battery capacity degradation using various sensor data, such as voltage, current, and temperature.
-RUL prediction is one of the critical functions of the PHM system, which forecasts the remaining cycles where the battery can be used safely.
-Accurate RUL prediction contributes to optimizing battery replacement timing and prevents accidents caused by excessive charging or discharging.
-RUL prediction methods using early cycling behavior can generally be classified into two approaches: model-based approach and data-driven approach [8, 9].
+SOH prediction is one of the critical functions of the PHM system, which forecasts the remaining cycles where the battery can be used safely.
+Accurate SOH prediction contributes to optimizing battery replacement timing and prevents accidents caused by excessive charging or discharging.
+SOH prediction methods using early cycling behavior can generally be classified into two approaches: model-based approach and data-driven approach [11, 12].
 However, model-based approaches often encounter challenges.
-In practice, the irreversible capacity fade is caused by an interaction of mechanical (e.g., particle cracking) and chemical (e,g., electrolyte decomposition) factors, making it quite difficult to mathematically identify these degradation mechanisms.
-Additionally, there can be a knee point where performance degradation, which progresses slowly in the early stages, accelerates after a certain cycle.
-This nonlinearity of performance degradation also contributes to the difficulty of model-based approaches.
-Machine Learning (ML)–based approach is a promising alternative to overcome this challenge [10, 11, 12].
-ML–based RUL prediction research typically employs time series algorithms, with changes observed for each charging/discharging cycle.
-Traditionally, Recurrent Neural Network (RNN) [13, 14, 15] and Long Short-Term Memory (LSTM) [16] have been used for such tasks: Liu et al.
+In practice, the irreversible capacity fade is caused by an interaction of mechanical (e.g., particle cracking) and chemical (e.g., electrolyte decomposition) factors, making it quite difficult to mathematically identify these degradation mechanisms.
+The existence of a knee point at which aging accelerates in a certain cycle, and often a slight capacity increase early in the cycle, makes model-based approaches difficult [13].
+Furthermore, while battery models such as the Doyle-Fuller-Newman (DFN) [14, 15] model have been successfully applied to predict SOH, the strong nonlinearity of these models necessitates further analysis, including identifiability analysis [16, 17].
+The data-driven approach is a promising alternative to overcome this challenge [18, 13, 19].
+The data-driven SOH prediction research typically employs time-series machine learning (ML) algorithms, with changes observed for each charging/discharging cycle.
+Normally, recurrent neural network (RNN) [20, 21, 22] and long short-term memory (LSTM) [23] have been used for such tasks: Liu et al.
 [24] propose an adaptive RNN that continuously updates the parameters of the network to improve the description performance of dynamic cycling behavior of the LIBs.
 Park et al.
-[18] conducted RUL forecasting using LSTM, but they struggle with learning long-term dependency [19].
+[25] conducted SOH forecasting using LSTM, but they struggle with learning long-term dependency [26].
 To overcome this limitation, the attention algorithm was introduced, enabling predictions that consider a broader context rather than just neighboring time steps [27, 28].
-Designed for natural language processing task and centered around the attention algorithm, Transformer architectures outperforms RNN and LSTM by taking a bird’s-eye view of time-series data to learn features of data, then trains in order of importance rather than sequential dependency [22].
+Designed for natural language processing tasks and centered around the attention algorithm, Transformer architectures outperform RNN and LSTM by taking a holistic view of time-series data to learn features of data, then trains in order of importance rather than sequential dependency [29].
 Chen et al.
-[23] conducted a prediction of RUL by Transformer, successfully overcoming the long-term dependency hurdle of RNN and LSTM.
-Despite being a game changer in time series forecasting, the Transformer has encountered difficulties when applied to time series prediction.
+[30] conducted a prediction of SOH by Transformer, successfully overcoming the long-term dependency hurdle of RNN and LSTM.
+Despite being a game-changer in time series forecasting, the Transformer has encountered difficulties when applied to time series prediction [31].
 Specifically, its high complexity in space (i.e., memory) time and its performance are not significantly superior to conventional LSTM models.
-An alternative, DLinear, was proposed to address these challenges [24].
-DLinear uses two linear layers considering trend and seasonality, in contrast to the complex structure and extensive computation of Transformer.
+To address these challenges, DLinear was proposed [31].
+DLinear uses two linear layers considering trend and seasonality, in contrast to the complex structure and extensive computation of the Transformer.
 Unlike the Transformer, DLinear takes both trend and seasonality into account in time series data predictions.
 Furthermore, it provides explainablity with its weights of linear layers, revealing parameter importance.
-For battery life, RUL typically follows a decreasing trend with occasional increases.
-While the Transformer struggles to learn this trend, DLinear captures both the decreasing and periodic increasing trends, outperforming LSTM and Transformer-based predictions in aspects of space complexity and time complexity.
-The performance of ML–based RUL prediction such as DLinear is determined by the quantity and quality of data.
+The SOH of the battery typically exhibits a decreasing pattern during its lifespan, with capacity increases occasionally observed at the early cycles, punctuated by periodic increases.
+Although the Transformer finds it challenging to grasp this trend, DLinear effectively captures both the downward and occasional upward movements, surpassing LSTM and Transformer-based predictions in terms of both space and time complexity.
+The performance of data-driven SOH prediction, such as DLinear, is determined by the quantity and quality of data.
 Since battery cycling experiments require a significant budget in time and cost, there are practical difficulties in generating large-scale datasets.
-Therefore, it is important to identify features that significantly contribute to battery lifespan and achieve high prediction performance even with a small number of training cycles.
+Therefore, it is essential to identify the key features that significantly contribute to battery health and to attain high predictive performance, even with a limited number of training cycles.
 However, because high-dimensional features can overfit the cycling behavior used for training, selecting key features is essential, and various feature engineering studies have been conducted to predict battery lifespan: Fei et al.
-[25] extracted 42 features from a large-scale lithium iron phosphate (LFP) battery dataset [11] and identified a high-performance prediction model structure by applying various feature selection approaches such as wrapper and filter approaches and various prediction models.
+[32] extracted 42 features from a large-scale lithium iron phosphate (LFP) battery dataset [13] and identified a high-performance prediction model structure by applying various feature selection approaches such as wrapper and filter approaches and various prediction algorithm.
 Hu et al.
-[26] proposed a fusion approach that combines the advantages of wrapper and filter approaches to select features that contribute to state-of-health prediction.
+[33] proposed a fusion approach that combines the advantages of wrapper and filter approaches to select features that contribute to SOH prediction.
 Greenbank et al.
@@
 [38], SHAP value-assisted feature selection was found to have superior performance compared to other algorithms for time series applications such as battery lifespan prediction.
-Moreover, the SHAP-based approach, which is a model-agnostic approach, is useful when applied to various machine learning models.
-In this study, we identify features that contribute to RUL prediction through the filter approach and SHAP-based approach and apply them to a ML model.
-Here, we identify the features that contribute to the battery lifespan and predict the RUL of the battery using the DLinear model.
-Features contributing to battery capacity degradation were extracted based on battery chemistry.
-For the entire feature set, each key feature subset identified through the filter approach and SHAP-based approach is used as input for LSTM, Transformer, and DLinear, and the accuracy and EoL cycle are compared.
-The rest of this paper is structured as follows.
-In Section 2, the three machine learning algorithms used for the prediction of RUL are described.
+Moreover, the SHAP-based approach, which is a model-agnostic approach, is useful when applied to various data-driven models.
+Based on the above considerations, this work focuses on feature engineering based on battery chemistry and data-driven prediction of SOH.
+Through sensitivity analysis, key features identified are used as inputs to a time-series forecasting model to accurately predict the decreasing trend of SOH.
+Specifically, our main contributions are presented below:
+  1. 1)
+    Using the data collected from the cycling experiment, a feature set required for SOH prediction is generated.
+PCC and SHAP-based global sensitivity analysis are performed to identify key features among all features.
+  2. 2)
+    The key feature subset is used as input to LSTM, Transformer, and DLinear, and the accuracy is compared with previous works.
+  3. 3)
+    The variability of cells composing modules or packs is intensively analyzed.
+Key features that are equally applied to the cells within specific modules or packs provide opportunities for efficient SOH prediction in practical BMS.
+This article is organized as follows: In Section 2, the three machine learning algorithms used for the prediction of SOH are described.
 In Section 3, the dataset used in the cycling experiment and the scheme for extracting features contributing to the battery lifespan based on battery chemistry are explained.
 Additionally, two algorithms used for feature selection are introduced.
-Section 4 discusses feature engineering and RUL prediction results, and Section 5 presents a summary of the results and future work.
-2 Time-series prediction model
-Figure 1: Schematic diagram of RUL prediction through feature engineering
+Section 4 discusses feature engineering and SOH prediction results, and Section 5 presents a summary of the results and future work.
+2 Time-series prediction algorithm
+Figure 1: Schematic diagram of SOH prediction through feature engineering
 2.1 Long short-term memory (LSTM)
@@
 RNN is the pioneering SNN structure designed for processing time-series data [20, 21, 22].
-The RNN model takes input to produce output, and also sends its output back to itself, to recurrent into the next time step.
+The RNN algorithm takes input to produce output, and also sends its output back to itself, to recurrent into the next time step.
 Due to this structure, each neuron in RNN has two sets of weights: one for input and another for multiplying with the output from the previous time step.
@@
 Furthermore, implemented with the Attention algorithm, the Transformer was devised.
-Recently, Transformer-based approaches and variant models of Transformer were utilized in predicting the RUL of LIBs [23].
+Recently, Transformer-based approaches and variant models of Transformer were utilized in predicting the SOH of LIBs [30].
 2.3 DLinear
@@
 Furthermore, due to its complex structure for analyzing sentences and texts written in languages, the Transformer required excessive computation compared to the conventional SNNs, thereby demanding a significant amount of time and memory.
-To address these two issues, DLinear was proposed [24].
-Algorithm 1 Decomposition-Linear (DLinear) Model for RUL Prediction
-1:Input: Feature matrix $X:[F_{1},F_{2},...,F_{20}]$, Target $y=\text{RUL}$
+The DLinear model was introduced to mitigate these two problems [31].
+Its inherently low computational burden, stemming from the linear structure of the model, offers a distinct advantage for BMS, as these systems necessitate real-time forecasting and are often constrained by the practical limitations of deploying models with excessive computational demands.
+Algorithm 1 Decomposition-Linear (DLinear) Algorithm for SOH Prediction
+1:Input: Feature matrix $X:[F_{1},F_{2},...,F_{20}]$, Target $y=\text{SoH}$
 2:Step 1: Compute the moving average $X_{t}$ using a sliding window
@@
 4:Step 3: Pass $X_{t}$ and $X_{s}$ through linear layers
-5:if $individual$ is True then
+5:if individualindividual is True then
 6:  for each feature $i$ in $X$ do
@@
 13:Output: Predicted time series data
-The detailed procedure for RUL prediction using DLinear is introduced in Algorithm 1.
+The detailed procedure for SOH prediction using DLinear is introduced in Algorithm 1.
 DLinear separates time-series data into Trend ($X_{t}$) and Remainder ($X_{s}$) using a moving average, and learns each through a separate linear layer.
@@
 3 Feature engineering
-In this section, feature engineering to extract parameters contributing to RUL from cycling data of LIBs and identify key parameters is explained.
-Our RUL prediction framework is validated using the NASA dataset [33].
-In Section 3.1, the process of extracting features expected to contribute to lifespan through information on voltage, temperature, and time for four LIBs cells (i.e., B0005, B0006, B0007, and B0018) is described.
+In this section, feature engineering to extract parameters contributing to SOH from cycling data of LIBs and identify key parameters is explained.
+Our SOH prediction framework is validated using the LIBs aging dataset provided by NASA’s Prognostics Center of Excellence [40].
+The nominal capacity of the four selected cells (i.e., B0005, B0006, B0007, and B0018) is 2 Ah.
+All the cells are cycled under the same charge and discharge procedure.
+In the charge phase, the cells are charged with a constant current (CC) of 1.5A until the voltage reaches 4.2V.
+After the voltage reaches 4.2V, they are charged with a constant voltege (CV) until the current reaches 0.02A.
+In the discharge phase, the cells are discharged in CC mode with 2.0 A until B0005, B0006, B0007, and B0018 reach 2.7V, 2.5V, 2.2V, and 2.5V, respectively.
+In Section 3.1, the process of extracting features expected to contribute to lifespan through information on voltage, temperature, and time for four lithium-ion batteries (LIBs) cells is described.
 In Section 3.2, the process by which a significantly contributing subset of the entire feature set is selected through the filter approach and SHAP-based approach is explained.
 3.1 Feature extraction
-In feature extraction, factors that contribute to RUL prediction are extracted based on battery cycling data collected from the Battery Management System (BMS).
+In feature extraction, factors that contribute to SOH prediction are extracted based on battery cycling data collected from the BMS.
 According to battery chemistry, the main factors that accelerate degradation are voltage [41] and temperature [42] during cycling, and degradation causes differences in charge and discharge profiles.
-Table 1 describes 20 features categorized by current and voltage, temperature, and time.
+Consequently, even if a feature that cannot be explained by battery chemistry is identified as a key feature [13], it is still necessary to find key features that can explain a battery’s health for SOH estimation.
+Table 1 describes 20 features generated from current, voltage, temperature, and time data collected in each cycle.
 Fig.
-2(a-c) shows that the current and voltage profile at the charging stage, temperature distribution, and discharge voltage curve are dependent on the RUL.
-These dependent characteristics are processed through statistical information such as skewness and included in the feature set.
+2(a-c) shows that the current and voltage profile at the charging stage, temperature distribution, and discharge voltage curve are dependent on the SOH.
+These dependent characteristics are processed through statistical information, such as skewness, and included in the feature set.
 In particular, the characteristic of the discharge curve, where differences occur after significant cycles, is explained by the slope at the early stage of the discharge (Fig.
@@
 Related parameter Feature designation Description
-Current and Voltage F1 Variance of measured voltage in discharging
-F2 Variance of measured current in discharging
+Current and Voltage F1 Variance of measured current in discharging
+F2 Variance of measured voltage in discharging
 F3 Median of loaded voltage in discharging
@@
 F20 Total discharging time
-Figure 2: Feature description of charging and discharging for B0005 battery cell: (a) CC charging time (F18), (b) Temperature in discharging, (c) Voltage in discharging, (d): Slope of discharge voltage curve (F6, F7, and F8)
+Figure 2: Feature description of charging and discharging for B0005 battery cell: (a) CC charging time (F18), (b) Temperature in discharging, (c) Voltage in discharging, (d) Slope of discharge voltage curve (F6, F7, and F8)
 3.2 Feature selection
@@
 The filter approach uses feature ranking to select key features.
-In general, the correlation between the extracted features and the objective function (i.e., RUL) scores the ranking of each feature, selects features that are higher than a certain threshold, and screens other features [36].
+In general, the correlation between the extracted features and the objective function (i.e., SOH) scores the ranking of each feature, selects features that are higher than a certain threshold, and screens other features [43].
 Because the filter approach is applied without being integrated into the prediction model, it can be easily applied to high-dimensional feature sets with relatively low computational budgets.
@@
 As for the key feature selection, all correlation coefficients between key features and cell capacity are evaluated based on Eq.
-3.
+(3).
 The three parameter subsets with the highest absolute correlation coefficients are regarded as the most sensitive features, while those with the lowest absolute values are considered the least sensitive features.
@@
 $\phi_{i}=\sum_{S\subseteq N\setminus\left\{i\right\}}\frac{\left|S\right|!\left(\left|N\right|-\left|S\right|-1\right)!}{\left|N\right|!}\left[f\left(S\cup\left\{i\right\}\right)-f\left(S\right)\right].$ (4)
-Where $\phi_{i}$ is the Shapley value of the $i$-th feature Fi, $N$ is the set of all features (e.g., in our case, $\left|N\right|$ is 20, see the Table 1), $N\setminus\left\{i\right\}$ is the set of all possible combinations of remaining features excluding feature Fi, $S$ is a subset of $N\setminus\left\{i\right\}$, $f\left(S\cup\left\{i\right\}\right)$ is the model prediction of the union of Fi and feature set $S$, and $f$($S$) is the model prediction for the features in $S$.
 4 Results and discussion
-Figure 3: PCC of all parameters and RUL for four battery cells: (a) B0005 cell, (b) B0006 cell, (c) B0007 cell, (d) B0018 cell
-Figure 4: SHAP score of all parameters and RUL for four battery cells: (a) B0005 cell, (b) B0006 cell, (c) B0007 cell, (d) B0018 cell
-The current, voltage, temperature, and time data collected from the cycling experiment were reproduced into 20 features as shown in Table 1.
+4.1 SOH prediction
+This work aims to provide insight into the SOH prediction research field by comparing widely used methods for processing data collected through BMS and generating prediction models for SOH prediction.
+In particular, many works are using PCC for feature engineering [45, 46, 47].
+In fact, we believe that global sensitivity analysis methods, such as SHAP and Sobol’ indices, can comprehensively analyze the influence of parameters, just like PCC.
+In addition to widely used time-series prediction models such as RNN, LSTM, and Transformer, advanced machine learning models were generated by integrating time-series prediction models with various machine learning algorithms, such as Convolutional Neural Network (CNN), and applied to SOH prediction (Table 2).
+Here, the DLinear, which has strengths in learning patterns of time-series data, is applied, and its accuracy is compared with that of LSTM and Transformer.
+According to Zeng et al.
+[31], the performance of the DLinear is far superior to that of the Transformer model in capturing temporal dynamic patterns between the set of continuous points.
+The SOH of a battery often shows a clear decreasing trend even when considering the capacity increase phenomenon in early cycles.
+Therefore, we believe that DLinear may be a suitable method for SOH prediction.
+However, for the fundamental performance comparison of the three algorithms, advanced models through integration with other ML algorithms are not considered.
+Table 2: Literature on advanced ML models for SOH prediction
+Reference Methodology Hyperparameter optimization
+Fu et al. [48] Variational mode decomposition(VMD)-Permutation entropy(PE)-Improved dung beetle optimization (IDBO)-Temporal convolutional network(TCN) IDBO
+Xia et al. [49] Bi-directional gated recurrent unit (BiGRU) Particle swarm optimization (PSO)
+He et al. [50] Gaussian process regression (GPR) Improved PSO with mutation factor and self-adaptive weight adjustment
+Zhou et al. [51] Convolutional bi-directional LSTM (CNN-Bi-LSTM) neural network Sparrow search algorithm (SSA)
+Zhang et al. [52] Mixed Kernel function relevance vector regression (MKRVR) Differential evolution (DE) with gray wolf optimizer (GWO)
+Zheng et al. [53] GRU-TCN-Support vector machine (SVR) (GRU-TCN: Meta learner, SVR: Base learner) Whale optimization algorithm (WOA)
+Tang et al. [54] CNN-BiGRU-Attention mechanism (AM) Kepler optimization algorithm (KOA)
+Ma et al. [55] LSTM Differential evolution GWO (DEGWO)
+Zhu et al. [56] BiLSTM
+Figure 3: SOH prediction results for four cells using all the features specified in the Table 1: (a) B0005, (b) B0006, (c) B0007, (d) B0018
 Fig.
-3 presents the results of PCC on 20 features for four cells.
-The three most sensitive features for each cell, as identified by the PCC, are:
+3 shows the prediction results of SOH for four cells using all the features specified in Table 1.
+In order to compare the SOH prediction accuracy for the four cells under the same criteria, we used up to the 70th cycle as training data.
+In addition, SOH of the four cells is defined as the initial capacity of each cell, not the nominal capacity.
+In this work, the EoL is defined as reaching 80% of the initial capacity, and the initial capacity of the four cells and EoL are different due to CtCV.
+DLinear shows considerable prediction results for all four cells, but LSTM and Transformer do not.
+In particular, DLinear shows lower root mean square error (RMSE) with fewer training cycles than previous works (Table 3).
+Despite the considerable prediction performance of the DLinear model, the prediction using all 20 features can often cause overfitting or increase the training time, which can be the main cause.
+Table 3: Comparison of RMSE across NASA battery cells with previous works
+Battery Ref. [25] Ref. [57] Our work
+Training cycles RMSE Training cycles RMSE Training cycles RMSE
+B0005 110 0.0168 80 0.0147 70 0.0055
+B0006 90 0.0152 80 0.0196 70 0.0055
+B0007 140 0.0085 80 0.0110 70 0.0084
+4.2 Feature engineering for SOH prediction
+Fig. 4 presents the results of PCC on 20 features for four cells. The three most sensitive features for each cell, as identified by the PCC, are:
   • B0005: F10 (Average temperature in discharging), F12 (Skewness of temperature in discharging), F13 (Minimum temperature in discharging)
@@
   • B0018: F7 (Slope of discharge voltage curve, 50s to 500s), F8 (Slope of discharge voltage curve, 50s to 1000s), F20 (Total discharging time)
-B0006, B0007, and B0018 include voltage characteristics during the discharging stage and time-related parameters during the charging or discharging stages as key parameters.
-In contrast, all three key parameters for B0005 are related to temperature characteristics, demonstrating significant variability across the four cells.
+In the B0005 cell, the three features (F10, F12, F13) with large values of PCC are all temperature-related features.
+On the other hand, the remaining three cells include voltage-related features in the discharge phase and time-related features in the charge or discharge phase as key features.
+The reason why the correlations with SOH are different, even though all four cells are cycled by the same charge and discharge protocol, can be explained by CtCV.
+In particular, it is noteworthy that no feature is identified as a key feature for all four cells.
+Figure 4: PCC of all features and SOH for four battery cells: (a) B0005 cell, (b) B0006 cell, (c) B0007 cell, (d) B0018 cell
+Figure 5: SHAP score of all parameters and SOH for four battery cells: (a) B0005 cell, (b) B0006 cell, (c) B0007 cell, (d) B0018 cell
+Fig. 5 shows the results of global sensitivity analysis based on SHAP for 20 features. The three most sensitive features for each cell, as identified by the SHAP, are:
+  • B0005: F2 (Variance of measured voltage in discharging), F10 (Average temperature in discharging), F11 (Variance of temperature in discharging)
+  • B0006, B0007, and B0018: F2 (Variance of measured voltage in discharging), F11 (Variance of temperature in discharging), F18 (CC charging time)
+It is noteworthy that B0006, B0007, and B0018 cells are all explained by the same features in the SHAP results.
+In addition, B0005 also includes F2 and F11 among the three features as key features.
+Unlike PCC, the results in SHAP show that there is a common feature subset across the four cells, which means that SHAP can provide feature selection results that successfully reflect CtCV.
+Figure 6: Comparison of results by time-series prediction model according to feature selection method and number of key features: (a) PCC-LSTM, (b) SHAP-LSTM, (c) PCC-Transformer, (d) SHAP-Transformer, (e) PCC-DLinear, and (f) SHAP-DLinear.
 Fig.
-4 shows the results of SHAP analysis, that F2 (Variance of measured voltage in discharging) and F11 (Variance of temperature in discharging) were selected as the highest values across all four cells.
-In addition to F2 and F11, F10 (Average temperature in discharging) was identified as one of the three most sensitive feature for B0005.
-Meanwhile, F18 (CC charging time), along with F2 and F11, was identified as one of the three most sensitive feature for B0006, B0007, and B0018.
-When training with the existing 70 cycles, the EoL prediction gap was well captured in Fig.
-5.
-EoL was defined as 80% of the nominal capacity.
-In the case of LSTM, it captured and tracked the decreasing trend well after the training cycles.
-However, as the prediction progressed toward the later stages, it failed to follow the downward trend, resulting in a flat graph.
-This is because of the gradient vanishing problem mentioned previously.
-The performance of Transformer was found to be inferior than LSTM due to the issue of over-stationarization, which is considered a drawback of the Transformer algorithm [38].
-Since the Transformer was originally developed for natural language processing rather than time-series data, it fails to capture critical temporal dependencies in non-stationary time-series data with clear downward trends.
-Despite the fact that forecasting non-stationary time-series data is crucial for real-world applications including battery RUL predictions, such over-stationarization limits the prediction capability of the Transformer in these non-stationary time-series data.
-Previous studies on RUL prediction using Transformers [23, 39] have demonstrated that Transformers outperform LSTM in terms of predictive performance with revised model of Transformer.
-However, this study focuses on a comprehensive analysis of feature engineering considering CtCV and the improvement in RUL prediction performance driven by the characteristics of the DLinear algorithm.
-In this context, optimization of the model architecture was not considered in this study.
-While LSTM and Transformer show a significant decrease in prediction performance as the training period decreases, DLinear maintained relatively accurate EoL prediction performance.
-Even when the training cycles were reduced to a maximum of 30 cycles, The performance of DLinear was observed to decline.
-Thus, DLinear demonstrated excellent performance in battery RUL prediction by separating the trend and seasonality of time-series data and learning them linearly, thereby effectively overcoming both the gradient vanishing problem of LSTM and the over-stationarization problem of the Transformer.
-Figure 5: RUL prediction of four LIBs : (a) B0005, (b) B0006, (c) B0007, (d): B0018
-Figure 6: Comparison of R2 score according to algorithm and training cycle
-$R^{2}=1-\frac{SS_{\text{res}}}{SS_{\text{tot}}}$ (5)
+6 shows the results applied to B0005 cell using 70 cycles for training to compare the accuracy of combinations of two feature selection methods and three time-series prediction models.
+We also investigated the effect of the number of features by training the model with three, four, and five features.
+This analysis aimed to validate our hypothesis that simply increasing the feature count does not guarantee better performance.
+For example, Fig.
+6(a) and Fig.
+6(d) present the results of training LSTM using features extracted from PCC and SHAP, respectively.
+Firstly, for LSTM, the results are almost similar between PCC and SHAP.
+Also, increasing the number of features does not necessarily increase prediction accuracy.
+A common phenomenon in all LSTM applications is that, after several cycles in the prediction phase, a flat SOH profile emerges, indicating that the LSTM model may encounter gradient vanishing issues.
+Secondly, for the Transformer model, PCC shows significantly improved predictive performance compared to LSTM.
+In the case of PCC, the prediction accuracy decreases substantially when four or five features are used, compared to using just three features.
+This suggests that including features with lower importance ranks may lead to a decrease in accuracy due to increased complexity rather than a model improvement.
+Furthermore, when using SHAP, it can be explained that the Transformer model does not effectively capture the patterns required for time-series prediction.
+[…]
+Figure 7: The SOH prediction of applying the key feature set of B0005 identified by PCC to other cells ((a): B0006, (b): B0007, (c): B0018) and (d) the SOH prediction of applying the key feature set of B0006 identified by SHAP to B0005
+4.3 Robust feature engineering to account for CtCV
+In practical BMS applications in EVs industry, battery modules or packs consist of dozens or hundreds of cells.
+Therefore, performing feature engineering individually for all these cells is practically impossible.
+Selecting a common set of key features for cells with the same battery chemistry will significantly contribute to practical SOH prediction.
 Fig.
-6 illustrates the R2 scores of the models based on each algorithm, while Table 2 presents the MSE values according to the algorithm and the size of the training cycles, and Table 3 shows the corresponding R2 scores.
-The R2 score is not affected by the scale of the data, therefore particularly useful for evaluating battery data with varying training cycle lengths and total cycle durations.
-In contrast, MSE, by squaring the errors, sensitively captures large errors in time-series data and is differentiable, making it a commonly used loss function in machine learning algorithms.
-When evaluated using MSE and R2 score, DLinear’s performance was outstanding.
-The bolded values in Table 2 and Table 3 indicate the superior RUL prediction performance of DLinear.
-Even when the training cycles were reduced to 40, DLinear achieved an R2 score of 0.995.
-$SS_{\text{res}}$ is the sum of squared differences between the observed values and the predicted values, and $SS_{\text{tot}}$ is the sum of squared differences between the observed values and the mean of the observed data.
-In cases where the model performs worse than simply predicting the mean of the observed values, the R2 score will be negative.
-This is because Eq.
-5 subtracts the ratio of the residual sum of squares from 1, meaning the resulting negative value is not constrained within the range of -1 to 1.
-Figure 7: RUL prediction of B0005 according to training cycle : (a) 70 cycle, (b) 50 cycle, (c) 40 cycle, (d): 30 cycle
-Figure 8: RUL prediction of B0006 according to training cycle : (a) 70 cycle, (b) 50 cycle, (c) 40 cycle, (d): 30 cycle
-Figure 9: RUL prediction of B0007 according to training cycle : (a) 70 cycle, (b) 50 cycle, (c) 40 cycle, (d): 30 cycle
-Figure 10: RUL prediction of B0018 according to training cycle : (a) 70 cycle, (b) 50 cycle, (c) 40 cycle, (d): 30 cycle
-Table 2: Comparison of MSE by algorithm according to training cycle
-Algorithm Training Cycle
-70 50 40 30
-B0005 LSTM 0.03407 0.01821 0.01738 0.01999
-Transformer 0.15731 0.31750 0.11027 0.37554
-DLinear 0.00003 0.00009 0.00019 0.00042
-B0006 LSTM 0.02038 0.00924 0.01089 0.01532
-Transformer 0.03699 0.24336 0.60980 0.35563
-DLinear 0.00003 0.00006 0.00009 0.00019
-B0007 LSTM 0.04070 0.01757 0.01530 0.01355
-Transformer 0.02553 0.19942 0.10073 0.15515
-DLinear 0.00007 0.00031 0.00062 0.00220
-B0018 LSTM 0.01101 0.03776 0.02583 0.01577
-Transformer 0.02165 0.02092 0.07549 0.20434
-DLinear 0.00002 0.00003 0.00003 0.00006
-Table 3: Comparison of R2 by algorithm according to training cycle
-Algorithm Training Cycle
-70 50 40 30
-B0005 LSTM -0.17398 0.66193 0.74823 0.75999
-Transformer -2.10452 -3.59821 -0.59773 -1.68900
-DLinear 0.98387 0.99705 0.99628 0.68536
-B0006 LSTM -0.37580 0.63463 0.69113 0.66219
-Transformer -1.49730 8.62568 16.29935 6.84430
-DLinear 0.99772 0.34021 -0.73222 -0.96263
-B0007 LSTM -0.69123 0.61148 0.74657 0.81998
-Transformer -0.06079 3.40917 -0.66853 1.06123
-DLinear 0.97061 0.99422 0.99511 0.33864
-B0018 LSTM -0.57772 -0.40026 0.38087 0.67313
-Transformer -2.10273 0.22419 -0.80949 3.21659
-DLinear 0.99418 0.98418 0.99464 0.95958
-In PHM technologies, the ultimate goal is to achieve accurate predictions with fewer training cycles as possible.
-To evaluate whether this study can contribute to PHM technology, we progressively reduced the training cycles to 70, 50, 40, and 30 and analyzed the prediction performance.
-Fig.
-7, 8, 9 and, 10 present the prediction results for four different battery cells (B0005, B0006, B0007, and B0018, respectively) while the number of training cycles was reduced.
-The corresponding MSE and R2 values are shown in Table 2 and Table 3.
-Reducing training cycles, LSTM exhibited poor performance, with its predictions deviating significantly from the actual values in the later stages, as shown in Fig.
-5.
-Similarly, due to the over-stationarization issue, predictions made using the Transformer also demonstrated severe fluctuations which was to minimize the loss function, consistent with the behavior observed in Fig.
-5.
-On the other hand, except for B0006 cell (Fig.
-8), DLinear demonstrates superior prediction performance compared to LSTM and Transformer even at 50 cycles and 40 cycles.
-However, except for B0018 (Fig.
-10), there is a considerable difference between the real cycling behavior and the predicted cycling behavior when the training cycles are set to 30.
-In fact, recent research using Transformer demonstrates superior predictive performance compared to LSTM [40, 23].
-However, this study focuses on feature engineering, and by using a simple network without considering improvements to the machine learning algorithms or hyperparameter tuning, it makes it easier to reproduce the results or apply this framework to other battery systems.
-SHAP and PCC were employed for feature selection.
-Although using the three most sensitive features of performance of PCC resulted in higher prediction performance, there was a lack of consistency across different batteries (B0005, B0006, B0007, B0018, respectively).
-In contrast, while the three most sensitive features identified using SHAP yielded slightly lower prediction performance than the performance of PCC, they demonstrated consistency in feature selection across different batteries.
-This result suggests that SHAP-based feature selection has a higher potential for transfer learning between batteries, suggesting a certain level of robustness.
-In battery capacity research, sampling is often conducted to ensure representativeness that accounts for differences between cells, underscoring the importance of generalization performance.
-The consistent selection of the same features across multiple batteries indicates that these features are likely useful in a general context rather than being specific to individual cells.
-This consistency also implies potential reductions in future data collection and analysis costs, making it significant for broader applications.
-Figure 11: RUL prediction of B0005 cell according to feature selection: (a) PCC, (b) SHAP
-Table 4: The most sensitive and least sensitive features in terms of importance, as determined by SHAP analysis
-B0005 B0006 B0007 B0018
-Three most sensitive 2a,*, 10, 11a 2a, 11a, 18a 2a, 11a, 18a 2a, 11a, 18a
-Three least sensitive 1, 7, 20 1, 5b, 14b 5b, 14b, 16 5b, 14b, 16
-* Based on the feature selection results using SHAP, the extracted features, denoted by superscript a, are identified as key features in three or more cells, while those, denoted by superscript b, are identified as not significant in three or more cells.
-Table 5: The three most sensitive and three least sensitive features in terms of importance, as determined by PCC
-B0005 B0006 B0007 B0018
-Three most sensitive 10, 13, 12 18, 20, 3 7, 8, 18 7, 8, 20
-Three least sensitive 17, 19, 20 13, 17, 15 13, 14, 15 10, 14, 16
-Fig.
-11 show the prediction results for B0005 cell using the sets of the three most sensitive features and the three least sensitive features, as analyzed through PCC and SHAP, respectively.
-The sets of the three most sensitive parameters and the three least sensitive parameters, as determined by SHAP and PCC analyses, are presented in Table 4 and Table 5.
-SHAP analysis reveals that more than half of the three most sensitive features are shared among batteries.
-A noteworthy point is, the key features identified by PCC differ significantly across the four cells, whereas those identified by SHAP analysis are identical for B0006, B0007, and B0018, wich B0005 sharing two features with the other cells.
-A possible interpretation of these results is that the key parameters identified through SHAP-based sensitivity analysis provide robust results concerning CtCV.
-In practice, the uncertainties propagated by various causes, such as differences in electrode thickness during manufacturing stages (e.g., Calendaring process), imply that cycling behavior can vary even when cycling the same battery system under identical conditions [5].
-The high robustness of SHAP suggests that it is a promising approach for predicting the RUL of large-capacity batteries (e.g., EVs) used in modules or packs.
-Meanwhile, as shown in Table 5, PCC revealed cases where a feature ranked low for one battery was ranked high for another.
-For example, F10, F13, and F20 were selected as the three most sensitive features based on PCC for some batteries, while they were in the three least sensitive features for others.
-These three features are related to temperature and time, and since batteries are significantly influenced by their surroundings, especially the ambient temperature, measurements such as maximum or minimum values can be affected by external factors or measurement errors.
-Therefore, these features appear to lack consistency in their significance across different batteries when evaluated based on PCC.
-Figure 12: RUL prediction of B0005 cell using two most sensitive features identified by SHAP: F2, F11
-In Table 4, F2 and F11 were consistently selected as the three most sensitive features across all four batteries.
-Building on this observation, we also examined the performance of predictions using only these two features.
-However, the performance did not match that of predictions made using three features.
-This suggests that, in addition to F2 and F11, including additional features such as F18, which encapsulate lifespan-related information, is recommended for improved prediction accuracy.
+7(a-c) shows the results of applying the key features (i.e., F2, F10, F11) identified by PCC on the B0005 cell to SOH prediction for B0006, B0007, and B0018 cell.
+The prediction accuracy decreases significantly when compared to the accuracy achieved by training with the unique set of key features identified by applying PCC to each individual cell (Fig.
+6(a-c)).
+On the other hand, when applying SHAP to identify the key features (i.e., F2, F11, F18) on B0006, B0007, and B0018 cell and then applying them to B0005 cell, the prediction accuracy is significantly higher (Fig.
+7(d)).
+One possible interpretation of these results is that using the same key features can yield acceptable performance for cells with the same battery chemistry.
+This implies that feature selection using SHAP offers robust feature selection results that can account for CtCV.
+Figure 8: Comparison of SOH predictions between feature selection methods according to training cycle: (a) 70 cycle, (b) 50 cycle, (c) 40 cycle, (d) 30 cycle.
+Despite the high robustness of the SHAP-based approach, the rapid decrease in prediction accuracy with reduced training cycles suggests the need for a feature selection method that maximizes both robustness and performance.
+When training for 70 cycles, all the cases using all features, PCC-based selection cases, and SHAP-based selection cases have fairly high prediction accuracies.
+When considering all features, there is no significant difference between the real SOH and predicted SOH even as the training cycles are reduced to 50, 40, and 30.
+However, when the number of training cycles decreases, the accuracy rapidly declines with SHAP.
+However, with PCC, predicted SOH remain very close to the real SOH with 50 cycles for training and still effectively track the trend of the real SOH when training cycles are reduced to 40 or 30.
+This suggests that PCC might be more appropriate for predicting the SOH of individual cells and during extremely early cycles, while SHAP could be more suitable for predictions at the module or pack scale.
 5 Conclusion
-PHM technology involves real-time diagnosis and prediction of the health state of a target system.
-Accurate RUL prediction helps implement condition-based management, where interventions are taken only when necessary by predicting the preemptive system replacement point before failures or defects occur.
-However, predicting failures through RUL requires measuring charge/discharge data during real usage, which is impractical and time-consuming during experiments.
-Therefore, it’s crucial to predict with higher accuracy using fewer experiments.
-In cases of clear trend-based predictions like battery degradation, DLinear outperforms models utilizing other algorithms.
-Even when trained on the same number of cycles, DLinear can predict further and more accurately than conventional LSTM or Transformer models.
-This suggests its potential contribution to implementing Condition Based Management for battery PHM technology in real industrial applications.
-On the other hand, identifying key features for predicting RUL aids in the implementation of PHM technology.
-PCC yields good performance for individual cells, but the key features vary significantly due to CtCV.
-Additionally, SHAP does not perform as well as PCC in terms of overall performance.
-However, SHAP consistently identifies key features, making it advantageous for battery research requiring generalization, despite the CtCV.
-Particularly, CtCV arises from both intra-cell factors such as morphology, architecture, and composition, and inter-cell factors like winding, connection, and assembly processes.
-Therefore, creating homogeneous batteries is challenging.
-In this regard, it is noteworthy that SHAP evaluates the importance of the same features despite CtCV.
+Accurate and efficient prediction of SOH is essential for realizing condition-based maintenance in EVs applications.
+In this work, we proposed an explainable SOH prediction framework combining feature engineering and the DLinear model.
+Leveraging public battery aging datasets, we demonstrated that DLinear outperforms conventional LSTM and Transformer models in both prediction accuracy and computational efficiency, even when trained with limited cycling data.
+To address the challenge of CtCV, we employed Pearson correlation and SHAP-based sensitivity analyses for feature selection.
+While PCC offered strong performance for individual cells, SHAP consistently identified key features applicable across multiple cells, enabling generalized SOH prediction for EVs battery modules.
+This is especially critical in large-scale battery systems, where individual cell-level modeling is impractical.
+The proposed framework offers a scalable, interpretable, and low-complexity solution with significant practical advantages for EVs BMS.
+The feature engineering process facilitates the generation of a dataset that is optimized for ML model training by enabling a substantial reduction in the number of features a BMS needs to measure and store.
+Furthermore, the use of SHAP ensures that key factors like CtCV are considered, while the DLinear model’s linear structure provides a substantially lower computational burden than more complex architectures like Transformer and LSTM.
+It enables early and reliable detection of battery degradation trends, contributing to safer operation, reduced maintenance costs, and improved battery utilization in electric vehicles.
+Future work will focus on integrating real-world EVs cycling data, refining model robustness under diverse usage conditions, and extending the framework for online learning in embedded BMS platforms.
 Code Availability
 The code used in this work will be available after publication.
+Appendix A Feature selection of B0005 cell
+Figure A-1: Two temperature parameters (F10 and F13) of B0005 cell
+Figure A-2: DLinear-based prediction results of B0005 cell using the two most sensitive parameters identified by SHAP